OpenStudy (anonymous):

Solve the equation 5/4x-x=2

OpenStudy (lgbasallote):

\[\large \frac{5}{4x} or \frac{5}{4}x\]

OpenStudy (anonymous):

Is it incorrect to multiply across by 4x?

OpenStudy (mimi_x3):

Or.. \[\large \frac{5}{4x-x} =2\]?

OpenStudy (lgbasallote):

depends...you might wanna tell us which is which first

OpenStudy (anonymous):

\[5\over4x \]

OpenStudy (anonymous):

Sorry for the confusion!

OpenStudy (mimi_x3):

\[\large \frac{5}{4x} -x=2\]??

OpenStudy (anonymous):

Yes

OpenStudy (mimi_x3):

Write it as: \[\frac{5}{4x}-\frac{x}{1} =2\] then times the top and bottom by \(4x\) for it to common.

OpenStudy (anonymous):

\[20x-4x^{2} = 8x\]

OpenStudy (mimi_x3):

Correct! Do you know what to do next? You can use the quadratic formula to solve for \(x\)

OpenStudy (anonymous):

Yeah, i got \[4x^2 + 20x - 8x = 0\] then divided across by 2. Minus B formula, it shouldnt be whole numbers right?

OpenStudy (anonymous):

X = 19.5 discard the negative value?

OpenStudy (mimi_x3):

I don't think that it's \(4x^2+2x-8x=0\) since when the equation was \(5-4x^2=8x\) you move the \(8x\) to the left..

OpenStudy (anonymous):

4x^2+20x−8x=0?

OpenStudy (anonymous):

if you multiply 5/4x by 4x, 4x cancels but you still multiply by the 5?

OpenStudy (mimi_x3):

It's \(4x^2+8x-5=\)

OpenStudy (anonymous):

AH! i made a fundamental mistake, :(

OpenStudy (mimi_x3):

\(4x^2+8x-5=0\)*

OpenStudy (anonymous):

Thank you mimi :)

OpenStudy (mimi_x3):

wait where did you get the \(20\) from?

OpenStudy (anonymous):

I thought you must multiply above & below the line in the fraction, or \[4x \times 5/4x\] so when the 4x cancelled, i multiplied the 5 by 4x! :(

OpenStudy (mimi_x3):

I think that I confused you; sorry. \[\frac{5}{4x}-\frac{x*4x}{1*4x} =2 =>\frac{5}{4x} -\frac{4x^{2}}{4x} =2 =>\frac{5-4x^{2}}{4x} =2\]

OpenStudy (mimi_x3):

Do you get it now?

OpenStudy (anonymous):

Yeah! Thanks so much for your time :)

OpenStudy (mimi_x3):

Np (: